Astronomy Report – December 7th

Astronomy Report

Name:   Akash Trivedi    Crew: 184
Date: 7/12/17

Sky Conditions: Clear

Wind Conditions: Light

Observation Start Time: 2:45pm

Observation End Time: 4pm

Summary: No interesting features, took a few pics

Objects Viewed: Sun

Problems Encountered: None

Journalist Report – December 6th

Crew 184 Journalist Report

Willie Schumann

06 December 2017

Title                            Climbing Higher

Narrative                  Day number five on Mars and we are still alive and happy to be here. While it continues to be quite chilly in the morning our water pipes didn’t freeze last night. Learning from yesterday we checked the pump first thing after waking up and while the hab tank filled up I looked around and saw into very content faces of relieved Marsonauts. The Generator is still a bit shaky, but our solar panels work great and as long as the sun is shining we should be fine.

We were planning to continue our Matryoshka project today on an EVA and we wanted to start quite early. That’s why we had an early lunch around ten o’clock. As I was assigned with commander Horn and Officers Trivedi and Hunt to join the exploration team I was also free of my daily cooking duties.

That gave me the opportunity to film First Officer Randazzo preparing delicious chicken noodle soup and mashed potatoes. This was particularly interesting for me, because it was a classic case of dried space food. We have milk powder, which we mix up with water. Just imagine the insane amount of milk we would have to fly up into space, matching our demand. It would probably smarter to ship a couple of space cows up to the red planet, but we would need their food… don’t get me started. This milk is essential for our breakfast cereal and it surely made our mashed potatoes creamy like on earth.

We really get our groove on with our EVA preparations. This time we pre-checked our radios and used the time efficiently in the pre-breathing chamber to get our suits and helmets on. Today I tried a new helmet, I had especially made on earth to meet my needs for filming and photographing, as it is important to have my analogue cameras close to my eye for focusing. It is a combination of a newly developed pressure resistant head and a 180° crystal face shield, which looks like a giant scuba mask. There were some issues with it during the EVA, but I will come back to that later.

So we drove out much quicker, then in the last EVA’s and took a rover and two ATV’s. Yesterday I made already some cool stabilized GoPro shots, heading into the Martian landscape. We repeated these shots today and I especially instructed the ATV driver to drive really close to the camera I had applied at the rear of our rover. It worked and they shots looked even more cinematic today.

For the geologists Matryoshka project of Science Officer Trivedi we climbed a few hills at different locations with interesting rock formations. The sun was up and we had some great frames with impressive shadows. My helmet turned a bit foggy, so a lot of my filming was a bit of a blind flight. But with the experience of the last days I managed quite well to deliver good shots. We stumbled about a stone formation, that almost looked like a bone of a giant creature. We could imagine it was a fossil of an early Martian life form. But maybe it is also our earthly expectation we project on Mars.

Being on these hills gave us a great overview on the territory we had covered in the first days we spent on the new planet. It was a quite majestic moment. Unfortunately one of my analogue cameras fell down twice today, opened up and also my settings were temporarily messed up. I will only know in a few weeks, if the pictures came out well. So the mission is to take further pictures, never to stop and to be sure to have enough material to show to the folks on the mother planet.

We came home a bit prematurely and were surprised to detect a little water leak just below our air lock chamber. We analyzed the scene and tried to analyze the source of the leak. Was it melted ice connected with the frozen pipes yesterday?     We later consulted mission control and were assured it is probably our kitchen sink, which shouldn’t bother us too much. And the Martian ground can hold a little bit more water for sure.

Back to the new helmet. Unfortunately it couldn’t be perfected for the Martian environment, just yet. At the moment breathing causes still moisture in the helmet, which makes it hard to see completely through the face shield. With the assistance of Health and Safety Officer Sczepaniak we applied some improvements to the breathing mechanism and first tests were very promising. I will test it in the field tomorrow. There are new perfect solutions, but I am happy, that my team is very good in troubleshooting. After all, a Mars mission is work in progress.

Yesterday night Commander Horn conducted a board game night, which was super fun. Crew Engineer Hunt showed unexpected winning potential, but I was happy with my performance as well. I did pretty well in an American trivia quiz, just my dice throwing skills leave much room for improvement. Evenings like these are very important for crew moral, which is still intact and promising for the days to come. Inline image 1

Personal Logbook             It was a good day. I am less exhausted today, even though it was equally tough on today’s EVA compared to yesterdays. Maybe I am getting used to the Martian life circumstances. Maybe I am evolving quicker than I thought.

I had to speak out on our EVA today to the rest of the crew. In lack of a crew camera for pictures of the geological surface I was constantly asked to take pictures of rocks. I can do that, but most of the time I am busy documenting every move of the crew. Doing these scientific shots can much easier be achieved by an additional camera of the respective specialist officer. Everybody agreed and I was happy, that my role is understood, and so I can continue my storytelling.

As I didn’t take a shaver with me I am growing an impressive beard at the moment. Maybe it will turn Hemingwayish beard in a few days. That would be a first for me, but very appropriate for an adventurer

Thank you very much for your help and attention.

Willie Schumann, Journalist, Crew 184

Sol Summary – December 6th

Crew 184 Daily Summary Report  06 Dec 2017

MDRS Daily Summary Report for Sol 5

Summary Title:  Continuing failures, schedule replans, and a successful EVA!

Mission Status:  We are still waiting for a nominal morning!  To fix yesterday’s bandwidth problems we powered off all devices in the HAB, we then connected our ‘internet’ laptop to the modem and watched 100MB of bandwidth disappear in ~ 5 minutes.  We concluded there was a background process of some sort operating on that computer as no known programs were running.  We were able to fix this issue by swapping out this laptop and powering down the fault one.  Communication restored!

We exulted in this victory for approximately 10 minutes, then noticed that our generator had failed again and we were back to a powerdown configuration.  Our SOC was holding at 47%, but it was again clear that we were operating with a degraded power system.  Communication with Mission Control has tentatively identified the problem as due to a ‘summer fuel’ blend which has been noted before, which is separate from the oil leak we observed earlier.  The generator will likely continue to be faulty for the next few days until we can procure a ‘winter’ blend.  Winters on Mars can be tough!.  We are continuing to conserve power wherever possible and using the generator minimally to supplement the solar arrays when required.

We resumed nominal operations this afternoon with two primary activities

We decided to finish up the EVA that we left incomplete yesterday, going to Matryoshka Site #2 which is adjacent to the MDRS HAB.  Our EVA prep procedure continues to be challenging, but we are getting better.  We have allocated the best radios and headsets for EVA (recently delivered!) and placed them adjacent to the EVA prep room.  We have also pre-staged all EVA equipment (boots / gloves / first aid kid / emergency water supply / maps / etc . . . ) in the EVA room to minimize the chance that our prebreathe is interrupted by having to go out to the HAB to get supplies.  We have also started carrying backup radios with us since we have had to cut an EVA short due to radios running out of batteries.  The EVA training continues!

Once outside we successfully found samples from numerous geological strata in our target era and bagged them for later delivery and chemical analysis.  We had some excess time upon returning so we also took some of the PR images requested and will uplink them as part of the journalist report.

When the majority of the crew got back to the habitat we did an observatory training sessions where the crew astronomy gave a tutorial in observatory and telescope operation.  There’s something about being on Mars that gives observatory operations extra gravitas as you are already in the ‘space’ mindset.

Sol Activity Summary:

1.      Matryoshka EVA Site #2

2.      Observatory Training Session (no observing)

3.      HAB Cleaning

4.      Reports and general HAB maintenance / checks

5.      Food inventory transfer to excel

6.      Exercise

Look Ahead Plan:

  1. Matryoshka EVA Site #3
  2. Nerve Block Training
  3. CBS News Visit (unconfirmed – request confirmation on whether media crew is coming from Mission Support)
  4. Filming and Interviews
  5. Solar Observation
  6. Communication Experiment (Time Delay)
  7. CPR Technique Training
  8. Exercise / General HAB Maintenance

Anomalies in work:

1.      HAB Leak from under EVA door (per Mission Control leak is expected and has been present for months, presumed to be from Kitchen piping)

2.      Generator failure

3.      Internet bandwidth problem was resolved today.

Weather: Still cold (below freezing), slightly windy

Crew Physical Status: All crew in good physical health.

EVA: See above for EVA details

Reports to be filed:

Greenhab

Operations

EVA Request

Journalist

Astronomy

Sol Summary

Support Requested:

1.      Trash pickup from rear airlock

2.      Replacement ATV Fuel

3.      Additional dish soap

4.  Confirmation of whether media visits will occur

GreenHab Report – December 6th

GreenHab Report

Trisha Randazzo

December 6, 2017

Environmental control: (Choose which is appropriate and explain further if needed)

Ambient (no heating or/cooling)

Ambient with window/door open

·         Below freezing for most of the day. Waited noon before opening door.

·         Temperature high 44, low 19

Heating

·         Functioning nominally, discovered heater comes on at 40 F

Cooling

·         Not available

Both heating and cooling

·         Nominal

Shade cloth on/off

·         On

Average temperature: (N/A until new sensor is delivered)

·         N/A

Changes to crops:  Note all emergence of seedlings, death of seedlings, etc

·         The same tomato is still doing poorly

·          Everyone else is growing quite well!!

·         Some new seedlings popped up in the two big garden beds

Daily water usage for crops:

·         Complete with no issues

Time(s) of watering for crops:

·         13:00

Morning research observations: (Currently not operational)

·         N/A

Changes to research plants:

·         N/A

Daily watering and amount of water used:

·         2 gallons used for all soils

Aquaponics:  (Currently not operational)

·         N/A

Narrative:  Any other information you want to share

·         N/A

Support/supplies needed:

·         N/A

Operations Report – December 6th

Crew 184 Operations Report  06/12/2017

 

SOL: 5

 

Name of person filing report: Josh Hunt

Non-nominal systems: Generator: Generator went off again overnight. Upon inspection of the generator, the fault light on the generator display was red-orange. Mission support was called and they were able to start up the generator again. Source of anomaly is believed to be the fuel itself which is still running on the “summer blend” and needs the “winter blend”. Winter blend has been ordered by Mission Support. Will log the time when the winter blend fuel arrives. Short-term forward plan is to turn off the generator at 10:00 pm tonight and turn beck on before 8:00 am tomorrow. 

                                   Hab Leak: EVA crew identified a water leak at the bottom of the Hab below the airlock porch. Mission support identified that this is a leak from the Hab Kitchen and is expected. Will monitor in case leak increases in severity in the future.

 

Generator— Hours run: Unknown (See Non Nominal Systems for details)

 

Solar— Morning SOC Charge was 47% 

             Noon SOC Charge was 100% 

             Evening SOC Charge (6:00pm) was 100%

 

Diesel – 46% Full

 

Propane –  72 Percent Volume

 

Gasoline (5 Gallon containers for ATV) –  2.5 Gallons.

 

Water (trailer) –  600 gallons

 

Water (static) –  400 gallons

 

Trailer to Static Pump used –  No

 

Water (loft) – Static to Loft Pump used –  Yes

 

Water Meter: 25 gallons

 

Toilet tank emptied: No

 

ATV’s Used: Yamahas 1 and 2

 

Oil Checked: Yes. All nominal.

 

Atv # Fuel Used Gals: Negligible

 

Tires Status: Yamaha 3 rear right tire is still wobbly. ATV is moved away for the time being and does not affect ops. Mission Control will address at their convenience.

 

# Hours the ATVs were Used today: 00:30 minutes

 

Deimos used: Yes

 

Hours: 00:30 minutes

 

Beginning charge: 87.4 %

 

Ending charge: 87.4 %

 

Currently charging: Yes.

 

 

 

HabCar used? No

 

Notes and Comments:

 

Summary of  Hab operations: Generator continues to provide trouble. This is the second night in a row that it has shut down ungracefully. Forward plan tonight is to check SOC at 10:00pm and then shut off the generator. 

 

 

Summary of GreenHab operations: All nominal

 

Summary of Science Dome operations: All nominal other than generator (see Summary of Hab Operations) 

 

Summary of health and safety issues: All nominal

 

 

Questions, concerns and requests to Mission Support: 

  1. Requesting extra fuel for ATVs. This is the same request previously submitted. We have 2.5 gallons of reserve fuel.
  2. Trash pickup from the rear airlock

Astronomy Report – December 6th

Name:  Thomas Horn

Crew:  184

Date:  12/6/2017

Sky Conditions:  Clear

Wind Conditions:  Light Wind

Observation Start Time:  3:30pm

Observation End Time:  5:00pm

Summary:  Crew Astronomer trained crew on observatory and telescope operations, no actual observing done.

Objects Viewed:  N/A

Problems Encountered:  None

Astronomy Report – December 5th

Name:  Thomas Horn

Crew:  184

Date:  12/5/2017

Sky Conditions:  Clear

Wind Conditions:  Still

Observation Start Time:  8:30am

Observation End Time:  9:30am

Summary:  Did a scan of solar features for unusual activity.  None
observed.  Image of solar surface attached.

Objects Viewed:  Sun

Problems Encountered:  None

Sol Summary – December 5th

Crew 184 Sol Summary Report  05 Dec 2017

MDRS Sol Summary Report for Sol 4

Summary Title:  Our first failure and major replan

Mission Status:  Today was a hectic day with three major failures our
team dealt with.
1.    Internet – Between 8:30am and 9:30am 800MB of bandwidth was
utilized. It is unknown what caused this but likely culprit was a
device being left on and doing a software update in the background.
Internet communication was lost at this time.  These reports are being
delivered by Shannon today.
2.    Water – During our morning fill of the loft tank the pump
was non-functional.  Possible causes were quickly narrowed down to
either a frozen pipe due to the below freezing temperatures or a
clogged filter. Our team immediately switched to using bottled water
which we keep as an emergency reserve while we troubleshot the issue.
At 1pm we called Mission Control for advice who recommended we connect
the pipe heating system and wait till midday, this resolved the issue
around 3pm and water was restored to the HAB
3.    Power – During our morning check of SOC it was noted that
the generator was not running, having failed sometime during the
night, and that SOC was at 3%.  We immediately instituted a shutdown
of MDRS powered equipment to preserve what little power we had and
investigated the generator.  It was noted that a significant oil leak
had developed from the oil cap, upon tightening the cap the leak
stopped and the generator was restarted.  When returning to the
science dome it was observed that SOC was hovering around 10% and did
not appear to be charging.  After consultation with Mission Control it
was discovered that the generator was in idle due to the power system
not requesting it to go to full delivered power given the fact that
the solar array system was charging the battiers (however slowly) .
In addition the SOC was likely incorrect due to it being a calculated
value and us confusing it due to going off scale.  We will assess SOC
tonight after a full day in the sun and decide forward plan.

Our morning was mostly consumed with maintenance and troubleshooting,
we quickly discovered that when receiving failures our level of
knowledge of the systems was exhausted in how to troubleshoot,
recommend future crew get access to and study manuals for MDRS
equipment prior to stay.

In the afternoon when we got our facilities in a stable configuratrion
four members of our team went on our first Matryoshka EVA. We were
able to go to the Kissing Camel ridge and collect two samples, however
we were unable to go to the other site in URC South due to delays in
getting out since we had to fix our non-nominal systems.

The two sample we collected were to highlight interesting
geomorphological features to be used in outreach activities back in
Oxford. The strike and dip of the ridge component the first sample was
extracted from was also taken. The intention of sample collection and
strike and dip measurement is to determine whether satellite and
outcrop scale imagery data is reliable at predicting these.

Sol Activity Summary:
1.    Troubleshooting and EVA as discussed above
2.    Exercise
3.    Assembling of exercise equipment using martian dirt
(medicine ball complete!)
4.    HAB Cleaning
5.    Solar observation

Look Ahead Plan:
With internet down we are unable to contact our scheduling team.  We
will receive the schedule in the morning when internet bandwidth
restarts and go over it as a team to plan our day.

Anomalies in work:
1.    Yamaha 250 ATV #3 wheel deemed non-recoverable.  We have
parked it off to the side for mission support.  Photos of damage
attached.

Weather: Very cold!  It was below freezing for most of the day.

Crew Physical Status: All in good health.

EVA: Delayed due to non-nominal systems, so we are requesting the same
EVA from Sol 3 but just for site A (mound on URC South) to finish off
the sample collection. The plan and crew utilised will be the exact
same as today.

Reports to be file:
Greenhab
Operations
Astronomy
EVA Request
Journalist
Sol Summary

Support Requested:
Trash placed in the Rear Airlock for pickup.  Burnable trash on the
side with the box.  Non-burnable trash on the other side.

Operations Report – December 5th

Crew 184 Operations Report 12/05/2017

SOL: SOL 4

Name of person filing report: Akash Trivedi

Non-nominal systems: Water (pump/lines – static to loft), generator
failure, sudden loss of data

Notes on non-nominal systems:  Water: tanks from static to loft not
transferring due to speculated frozen pipe or blockage. Waited until
10am as per protocol to see if frozen pipes thawed. At 10am, still
frozen. Waited until 12pm for sunlight to be in contact with pipes to
see if thaws. After the meeting at lunch, we learned that a plug by
the static tank needed to be seated. This allows heating along the
pipes. Once plugged we waited some time, and pump was operational
again at around 3pm. Anomaly resolved.

Generator: Noticed generator off at 0900 having failed during the
night. Turned on and seemed to be operating nominally. Returned to
check up on it at 1000, and noticed a coolant leak from the radiator.
Generator turned off and leak corrected by tightening the antifreeze
cap. Generator turned back on and checked status on power console –
seems to not be doing the charging as it’s on “absorb” and SOC not
increasing despite minimal hab power usage. After Lunch meeting, we
monitored thegenerator closely and at sunset the SOC returned to 100%.
Will continue to monitor closely, but power is good at this time.

Loss of internet data: Data check was conducted at 0845 when it was
approximately 800MB, and an hour later, this dropped to zero. Reason
unknown.

Generator (hours run):  Unknown due to night time failure

Solar— SOC % (Before generator is run at night) 4% at 0900, 14% at
1000 (when coolant leak discovered), static 9% on “absorb charging”
with minimal hab usage and generator powered on

Diesel – 48%

Propane –  79 percent volume

Ethanol Free Gasoline (5 Gallon containers for ATV) – 1.25 gallons

Water (trailer) – 600 gallons

Water (static) – 450 gallons

Trailer to Static Pump used – No

Water (loft) – Static to Loft Pump used – Yes

Water Meter: 46 gallons

Toilet tank emptied: Yes

ATV’s Used: Honda ATV, Yamaha 1

Oil Added? No

ATV Fuel Used: 1.25 gallons

# Hours the ATVs were Used today: 00:45

Notes on ATVs: Yamaha 3 Rear right tire is still wobbly, EVA team
assessed and determined it will require. Photos will be provided in
overall Sol 4 package for awareness and review.

Deimos rover used: Used

Hours: 00:45

Beginning charge: 100%

Ending charge: 100%

Currently charging:Yes

Sojourner rover used:  ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used:  Not used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used:  Not Used

Hours:

Beginning charge:

Ending charge:

Currently charging:

HabCar used and why, where?  No

General notes and comments: Multiple non-nominal systems!!!

Summary of internet: Sudden anomalous data loss – see above

Summary of suits and radios: All nominal based on last usage

Summary of Hab operations: Loft water tank almost empty, multiple
non-nominal systems

Summary of GreenHab operations:  All nominal

Summary of ScienceDome operations: All nominal

Summary of RAMM operations: Not Operational

Summary of health and safety issues: No problems now that water, heat,
and power are back.

Questions, concerns and requests to Mission Support: See above on
non-nominal systems. Lack of training and relevant manuals for
troubleshooting significant issues in mission critical systems. Poor
insulation of static tank to loft tank pipes – could lead to pipe
cracking